OpenStudy (anonymous):

help with graphing this Parabola: x = y^2 - 9

OpenStudy (anonymous):

Put these two halves together in the images below. Basically it's a parabola like normal only turned sideways because the independent and dependent variables have been switched. \[x=y^2-9\] \[x+9=y^2 \] \[\pm \sqrt{x+9} = y \]

OpenStudy (anonymous):

okay thx...but is misread the question given and they wanted the quadrants, which I know lol im sorry i wasted your time XD :)

OpenStudy (anonymous):

This was still very helpful though!!

OpenStudy (anonymous):

Plug in when x=0 to find the y-axis (vertical) intercepts, you should get two answers, which you'll see on the graph. (hint: What's the square root of 9?) Ah no worries, do I get a medal for answering the wrong question correctly then? haha

OpenStudy (anonymous):

^_^ This is a good tip for all parabolas, knowing how to recognize the ones going "sideways"

OpenStudy (anonymous):

YEA I GAVE YOU A METAL LOL!!